# Improving Local Training in Federated Learning via Temperature Scaling

### v2

@@ 3.2 FLex&\&&Chill and Logit Chilling
The lower temperatures used by Logit Chilling directly influence the local training process of each federated learning client. Notably, the seamless integration of Logit Chilling and FLex&\&&Chill with existing federated learning frameworks is a distinct advantage, allowing for a straightforward implementation within already deployed frameworks. Furthermore, we note that FLex&\&&Chill can be used orthogonality with other efforts that accelerate federated learning training operations such as FedProx and SCAFFOLD. For a detailed explanation and experimental result about the orthogonality of FLex&\&&Chill, please refer to Appendix  C.
4 Experiment
@@
We evaluate the performance of FLex&\&&Chill using three datasets: (i) FEMNIST from the LEAF database Caldas et al. (2018a), including the Extended MNIST handwriting data Cohen et al. (2017) , (ii) the CIFAR10 dataset, and (iii) the CIFAR100 dataset Krizhevsky et al. (2009). As metrics, we use the average accuracy and training loss observed at the clients, selected to observe both the performance gain and the effectiveness of FLex&\&&Chill in its training operations. Furthermore, we also present the number of federated learning rounds needed to achieve a pre-defined target accuracy.
(a) FEMNIST-DNN
(b) CIFAR10-CNN
(c) CIFAR100-ResNet18
Figure 4: Average federated learning client-side model accuracy for different temperatures used in training. Overall, lower temperatures show higher average model accuracy for federated learning systems.
@@ 4.1 Experiment Setup
Dataset. We detail the datasets used in our work as follows.
@@
  • FEMNIST dataset: The FEMNIST dataset comprises of 10 handwritten digits (0-9) and 52 characters (26 lowercase and 26 uppercase) images contributed by 712 users, totaling 157,132 samples. We selected a subset of 36 users, each with diverse amounts of local data to account for the non-i.i.d. environment. We followed the implementations provided by the original work Caldas et al. (2018b).
  • CIFAR10 and CIFAR100 dataset: We use baseline CIFAR10 and CIFAR100 datasets Krizhevsky et al. (2009) and manually partition data over 50 federated learning users, with each holding 1,000 samples.
Model. For each dataset, we utilized distinct model architectures, taking into account the specific challenges and characteristics inherent to each dataset. For FEMNIST, an MLP model featuring four linear layers was employed. We term this combination as “FEMNIST-DNN” in the rest of our evaluations. In the case of the CIFAR10 and CIFAR100 datasets, experiments were conducted using a two-layer CNN model McMahan et al. (2017) and a ResNet18 He et al. (2016), respectively. These two cases are denoted as “CIFAR10-CNN” and “CIFAR100-ResNet18.”
Experiment Details. Each case underwent local training for 300 federated training rounds, utilizing FedAvg McMahan et al. (2017) as the baseline federated learning framework. FedAvg is a widely used framework for federated learning. In each round, the server selects 10 participating clients, employing Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.001 Ruder (2016). Each client trained its model for 10 local epochs, employing a batch size of 16, and utilized cross-entropy loss as the loss function. Unless specified otherwise, the experiment was conducted using the same hyperparameters as mentioned in this section.
4.2 Average Client-side Accuracy
We report the average client-side model accuracy obtained for models trained with different temperatures in Figure 4. For all three datasets (and the three respective models used for each test case) we see an increasing accuracy trend observed for lower temperatures. Note that for the FEMNIST-DNN case with $T$=0.05, we noticed that the accuracy converges after 120 rounds; thus, perform an early exit for this case. One interesting aspect we can notice is that the accuracy for the $T$=0.05 case, which is the lowest temperature we tested for, does not show the best performance for the CIFAR100-ResNet18 experiment. While the $T$=0.05 case shows the best performance of FEMNIST-DNN and CIFAR10-CNN, for CIFAR100-ResNet18, we see an accuracy worse than the $T$=0.5 case. This suggests that training with the lowest possible temperature value is not always ideal, rather this result highlights the importance of identifying a proper training temperature as we will further discuss in Section 5. A closer zoom into the plots suggests that the best-performing configuration can change over the training round and can vary for different datasets, but overall, we can notice that exploiting a temperature value lower than “1” in the training process can help improve the classification accuracy.
The average accuracy for all federated learning clients after 300 training rounds presented in Table 1 confirm our observations, indicating that with $T$=0.05, the CIFAR10 case shows a 3.37% improvement in accuracy compared to the $T$=1 case. Even for other cases, the use of a low-temperature value shows promising results in terms of accuracy. We emphasize once more that this does not mean that the accuracy will always perform better at the lowest temperature. Rather, our results suggest the need to carefully explore the fractional temperature space in the federated learning training process. We note that the performance for model performance and convergence speed tends to be similar regardless of federated learning hyperparameters (see Appendix D).
4.3 Federated Learning Convergence Analysis
Table 1: Average classification accuracy, model training loss, and number of federated learning training rounds needed to achieve the maximum accuracy of the $T$=4 cases for each dataset-model combination.
Finally, in the last columns for each test case in Table 1 we present the number of federated learning training rounds needed to achieve a preset target accuracy for different temperature values used for model training. This number provides an understanding of how efficient the local training processes are when different temperature values are applied. Specifically, in this experiment, we configure the target accuracy to be the maximum accuracy achieved within 300 training rounds with $T$=4. For example, for the FEMNIST-DNN and CIFAR10-CNN cases, the maximum accuracy for $T$=4 was achieved at the 299th round, while CIFAR100-ResNet18 achieved such accuracy after 288 rounds.
From the results in Table 1 we can notice that with $T$=0.05, the FEMNIST-DNN case achieves the target accuracy after only 22 training rounds. We see similar trends of low temperatures positively affecting the training efficiency throughout our results. This is from the general aggressiveness of training behaviors when applying lower temperature training. Given a single training sample, while high temperature-based training conservatively absorbs the available information, lower temperatures are more progressive in making model changes. Given the small datasets used for local training and the need to effectively deliver its knowledge to the server for aggregation, we see FLex&\&&Chill as an effective way to train client-side models in the federated learning process.
4.4 Feature Space Similarity Analysis
(a) $T$=0.05
(b) $T$=1.0
Figure 6: Client-wise feature space similarity matrix measured in penultimate layers by using CKA and inference accuracy in different temperatures (0.05, 1.0) and federated rounds (1, 5, 50, 100). (INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
To better understand the mechanism and rationale behind the performance of FLex&\&&Chill, we analyze the similarity of the feature space between global (aggregated) models trained with different temperatures against a reference model trained in an i.i.d. data configuration. Specifically, we employ the Centered Kernel Alignment (CKA) Kornblith et al. (2019) for computing the similarity of feature spaces, a widely used approach in evaluating model divergence in federated learning Collins et al. (2022); Luo et al. (2021). Note that a limitation of CKA measures is that they only report network similarity without considering training quality. Thus, we use a well-trained reference model from an i.i.d. environment to observe CKA variations of the penultimate and classification layers between the global models and the reference model.
We note that the reference model is trained with $T$=1, while comparison models (e.g., global models from federated learning) are trained in non-i.i.d. environments with different $T$. In the experiments, we used a 2-layered CNN and the CIFAR-10 dataset as in Section 4.1, with the dataset distributed among 10 clients using a Dirichlet distribution Hsu et al. (2019). All clients participated in the federated training process. For detailed discussions on the experimental setup and CKA, please refer to Appendix E.
As the accuracy and CKA plots for different layers in Figure 5 show, introducing the non-i.i.d. environment itself drops the model accuracy (73.86% $\xrightarrow{}$ 58.19%) at $T$=1, with the CKA of the penultimate/classification layers between the two models remaining at 0.7644/0.8214 after 300 rounds. Nevertheless, with $T$=0.05, despite the dataset disparity, the global model achieved a comparable accuracy of 72.76% with relatively higher CKA values (0.8898/0.9026) in both the penultimate and classification layers. We point out that improved gradient flows via applying low temperatures help models to learn similar local features (i.e., layers 1, 2), affecting the following layers (i.e., layers 3, 4) in the early stage of the learning process. As a result, applying low temperatures during local training effectively improves the overall accuracy performance in a non-i.i.d. data distribution setting and helps models learn similar feature spaces compared to the ideal i.i.d. data configuration.
We now target to better understand how local model updates progress with different $T$. For brevity of the paper we only report the CKA at the penultimate layer, and plot all other cases in Appendix E. Specifically, Figure 6 plots the model accuracy (ACC) and measured CKA at the penultimate layer of a client’s local model with (i) the initially distributed model (INIT), (ii) the aggregated model (AGG), (iii) the ideal reference model trained with i.i.d data distribution (REF), and (iv) other client models while $T$=0.05 and 1.0. In Round 1, clients’ local models trained with $T$=1 show high feature space similarity against each other while failing to learn the features of the ideal reference model. With $T$=0.05, the CKA between local models are relatively low while showing high similarity with the reference model. After five rounds (and beyond), local models trained with both $T$=1.0 and $T$=0.05 showed high similarity amongst themselves and the initial/aggregated models, but the $T$=0.05 case continuously shows higher similarity with REF and reports higher accuracy. Hence, we argue that adapting low temperatures allows local models to learn effective and diverse features to quickly converge the federated learning process.
4.5 Impact of Dataset Dispersion and Federated Learning on FLex&\&&Chill
We now examine the impact of FLex&\&&Chill on different model training configurations. Specifically, using the CIFAR10-CNN test case, we examine two additional configurations: (i) centralized learning with i.i.d. data and (ii) federated learning with i.i.d. data. Note that the federated learning network is identical to our previous experiments except for the dataset distribution and was tested for 300 rounds. By comparing these results with the non-i.i.d. federated learning results, we target to understand the impact of FLex&\&&Chill on federated learning and dataset disparity in greater detail.
As Table 2 shows, applying FLex&\&&Chill, thus a lower $T$ does not induce a positive impact on the performance of centralized learning scenarios. In fact, for centralized learning, controlling $T$ for the training process does not show a noticeable impact. We see the same phenomena for the federated learning system with i.i.d. samples as well, suggesting that the performance enhancements we see in Table 1 are not an effect of the federated learning process itself.
Table 2: Average inference accuracy and model training loss for CIFAR10-CNN with different configurations. The impact of FLex&\&&Chill is prominent for federated learning scenarios with non-i.i.d. datasets, whereas its impact is less than elsewise.
We do point out that holding i.i.d. datasets in federated learning will show overall higher performance compared to non-i.i.d. configurations (regardless from $T$), but is not a practical case to assume when deploying federated learning systems. Rather, as we discussed, the impact of exploiting low training temperatures is beneficial when dealing with non-i.i.d. datasets.
As discussed in Section 3, low temperatures expedite local training even with a small amount of training data. Nevertheless, the effect of low-temperature training on the model generalization and aggregation stability of federated learning has not yet been thoroughly studied. To inspect the stability of federated learning aggregation for models trained with low temperatures, we measured the performance change of each client-side model before and after the server-side aggregation round.
5 Discussion
$\bullet$ Identifying the optimal temperature. Our study exhibits the effectiveness of exploiting lower temperatures as a way to enhance federated learning performance. However, since the softmax layer training temperature can be considered a hyperparameter, optimal temperatures can vary depending on the applications and model architectures. Thus, like other hyperparameters used in model training (e.g., batch size, local epoch count), identifying an optimal temperature value is challenging to analytically solve, being beyond the scope of this paper.
$\bullet$ Dynamic per-sample temperature scaling. While our paper focuses on understanding the advantages of maintaining a constant low temperature for the training process, this may not be the ideal approach when targeting maximum performance. In reality, varying temperatures can be applied in the training process with respect to their significance Shin et al. (2022); Zeng et al. (2021). That would mean that the training process can exploit lower temperatures for the samples that are “harder” to learn, while higher temperatures are used for those easier to understand. Alternatively, a more course-grain approach of dynamically adjusting the temperature on a per-federated learning round-basis can be another meaningful approach to adapt towards changing local data characteristics.
$\bullet$ Supporting heterogeneous federated learning. While many federated learning frameworks assume homogeneous devices, recent work points out that this is not always the case Park and Ko (2023). In a federated learning network consisting of devices with heterogeneous computing/network capabilities, the issue of “stragglers”, which take longer to complete local training, becomes important to address. Leveraging insights gleaned from our empirical observations, applying lower temperatures for potential stranglers can be an interesting way of uniforming the local training duration. This would require a prior analysis on the conditions expected at each device and a global consensus on the temperature configuration. We find this as an interesting topic that requires further investigation.
6 Conclusion
This research introduces a novel training approach for federated learning systems, FLex&\&&Chill, aimed at enhancing convergence and improving model performance by exploiting a temperature scaling method termed logit chilling in federated learning local model training operations. Specifically, our work revolves around the implications of employing fractional temperature values ($T\in(0,1)$) during the model training process. Our findings, derived from an extensive set of evaluations using three datasets and three widely applied neural networks for federated learning, demonstrate that exploiting the FLex&\&&Chill approach in the client-side model training process holds the potential to expedite model convergence by inducing more pronounced changes despite a limited number of local samples. Consequently, we show that by doing so, we can also achieve an overall higher inference accuracy. While the impact of lower temperature varies with respect to the neural network architecture and dataset, our evaluations reveal up to 6.00$\times$ improvement in federated learning model convergence speed and 3.37% improvement in inference accuracy. We see this work as a first step into devising a more fine-tuned (and optimal) training mechanism for federated learning systems.
Appendix A Gradient of softmax with varying temperature
Proof.
Let cross-entropy loss for softmax as $\mathcal{L}=-\sum_{j=1}^{C}y_{j}\log{(p_{j})}$, where $C$ is the number of classes and $y_{j}$ and $p_{j}$ corresponds to the hard label (i.e., onehot-encoded label) and probability for given class. Here, the probability $p_{j}$ is computed by exploiting the softmax with varying temperatures which can be formalized as follows.
$p_{j}=\frac{e^{z_{j}/T}}{\sum_{i=1}^{C}{e^{z_{i}/T}}}$
$\begin{split}\frac{\partial\mathcal{L}}{\partial z_{i}}&=-\sum_{j=1}^{C}\frac{\partial y_{j}\log{(p_{j})}}{\partial z_{i}}\\ &=\sum_{j=1}^{C}y_{j}\frac{\partial\log{(p_{j})}}{\partial z_{i}}\\ &=-\sum_{j=1}^{C}y_{j}\frac{\partial\log{(p_{j})}}{\partial z_{i}}\\ &=-\sum_{j=1}^{C}y_{j}\frac{1}{p_{j}}\frac{\partial p_{j}}{\partial z_{i}}\\ &=-\frac{y_{i}}{p_{i}}\frac{\partial p_{i}}{\partial z_{i}}-\sum_{j\neq i}^{C}\frac{y_{j}}{p_{j}}\frac{\partial p_{j}}{\partial z_{i}}\\ &=-\frac{y_{i}}{p_{i}}\frac{1}{T}p_{i}(1-p_{i})-\sum_{j\neq i}^{C}\frac{y_{j}}{p_{j}}(-p_{j}p_{i})\\ &=-\frac{y_{i}}{T}+\frac{1}{T}y_{i}p_{i}+\frac{1}{T}\sum_{j\neq i}^{C}y_{j}p_{i}\\ &=-\frac{y_{i}}{T}+\frac{1}{T}\sum_{j=1}^{C}y_{j}p_{i}\\ &=-\frac{y_{i}}{T}+\frac{1}{T}p_{i}\sum_{j=1}^{C}y_{j}\\ &=\frac{1}{T}(p_{i}-y_{i})\\ \therefore\frac{\partial\mathcal{L}}{\partial z_{i}}&=\frac{1}{T}(p_{i}-y_{i})\\ &=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})\end{split}$
Appendix B Experimental Details
B.1 Model Architecture
Layer Details Repetition FC 1 Linear(784, 512) ReLU() ×$\times 1$ FC 2 Linear(512, 256) ReLU() ×$\times 1$ FC 3 Linear(256, 128) ReLU() ×$\times 1$ FC 4 Linear(256, 62) ×$\times 1$
Layer Details Repetition Conv 1 Conv2d(3, 10, k=(5, 5), s=(1, 1)) ReLU() MaxPool2d(k=(2, 2)) ×$\times 1$ Conv 2 Conv2d(10, 20, k=(5, 5),s=(1, 1)) ReLU() MaxPool2d(k=(2, 2)) ×$\times 1$ FC 1 Linear(500, 256) ReLU() ×$\times 1$ FC 2 Linear(256, 10) ×$\times 1$
For the ResNet18 model used to train the CIFAR100 dataset, we exploited PyTorch implementation Paszke et al. [2019] by only replacing the final classification head from Linear(512, 1000) to Linear(512, 100).
B.2 Hyperparameters
Hyperparameters FEMNIST-DNN CIFAR10-CNN CIFAR100-ResNet18 CKA Evaluation total clients 36 50 10 communication rounds 300 100 optimizer SGD learning rate 0.001 0.005 local epoch 10 10 participants per round 10 10 batch size 16 32 shard size | $\alpha$ N/A 200 $\alpha$=0.5 loss function Cross Entropy
B.3 Data distribution
@@
B.4 Hardware
For the experiments, we exploited server with four GeForce RTX 2080 Ti GPUs, Intel(R) Xeon(R) Silver 4210 CPU @ 2.20GHz, and 64GB RAM.
Appendix C Orthogonality of FLex&\&&Chill
C.1 FLex&\&&Chill As An Algorithm
We formulate the operations of FLex&\&&Chill using Algorithm 1. Here, typical federated learning operations take place at both the server and clients except for the fact that the server determines a target $T$ at each round and local model training operations are performed with $T$. Since FLex&\&&Chill only manipulates the local training process which inherently provides orthogonality to our proposed method, it is extremely easy to integrate with existing federated learning frameworks.
Algorithm 1 FLex&\&&Chill
Data ($D_{1},D_{2},...,D_{N}$) where $D_{i}$ is the user $i$’s local data.
Model ($w^{1}_{t},w^{2}_{t},...,w^{N}_{t}$) where $w_{t}^{i}$ is the user $i$’s model in $t$-th federated round.
Parameter: Temperature $T\in(0,1)$
1:Server executes:
2:Initialize $w_{0}$
3:for each round $t=1,2,...,R$ do
4:    $S_{t}\xleftarrow{}$ (random set of $m$ clients)
5:    for $k\in S_{t}$ do
6:         $w_{t}^{k}\xleftarrow{}$ ClientUpdate($w_{t},T$)
7:    end for
8:    for $k\in S_{t}$ do
9:         $w_{t+1}\xleftarrow{}$ Aggregation($w_{t}^{k}$)
10:    end for
11:end for
12:ClientUpdate($w_{t}$, $T$):
13:$w_{t}^{k}\xleftarrow{}w_{t}$
14:for data, label $\in D_{k}$ do
15:    $\mathbf{z}=w_{t}^{k}(data)/T$
16:    $l(\mathbf{z})=loss(\mathbf{z};label)$
17:    $w_{t}^{k}\xleftarrow{}LocalUpdate(w_{t}^{k},l(\mathbf{z}))$
18:end for
19:Return $w_{t}^{k}$ to server
C.2 FLex&\&&Chill with Advanced FL Schemes
As mentioned in Section 3.2, FLex&\&&Chill is an orthogonal approach to enhance the federated learning process. To demonstrate this, we trained the 2-layer CNN with FedProx Li et al. [2020] and SCAFFOLD Karimireddy et al. [2020] representing more recently proposed FL schemes compared to FedAvg used in this work.

### v3

@@ 3.4 FLex&\&Chill and Logit Chilling
The low temperatures used by Logit Chilling directly influence the local training process of each federated learning client. Notably, the seamless integration of Logit Chilling and FLex&\&Chill with existing federated learning frameworks is a distinct advantage, allowing for a straightforward implementation within already deployed frameworks. Furthermore, as mentioned, we note that FLex&\&Chill can be used orthogonally with other efforts that accelerate federated learning training operations, such as FedProx, SCAFFOLD, and FedBN. We provide an open-sourced implementation of FLex&\&Chill at https://github.com/eis-lab/temperature-scaling.
4 EVALUATION
@@
We evaluate the performance of FLex&\&Chill using three datasets: (i) FEMNIST from the LEAF database [caldas2018leaf], including the Extended MNIST handwriting data [cohen2017emnist], (ii) the CIFAR10 dataset, and (iii) the CIFAR100 dataset [krizhevsky2009learning]. As metrics, we use the average accuracy and training loss observed at the clients, selected to observe both the performance gain and the effectiveness of FLex&\&Chill in its training operations. Furthermore, we also present the number of federated learning rounds needed to achieve a pre-defined target accuracy.
@@ 4.1 Experiment Setup
Dataset. We detail the datasets used in our work as follows. (i) The FEMNIST dataset comprises of 10 handwritten digits (0-9) and 52 characters (26 lowercase and 26 uppercase) images contributed by 712 users, totaling 157,132 samples. We selected a subset of 36 users, each with diverse amounts of local data to account for the non-i.i.d. environment. We followed the implementations provided by the original work [caldas2018leaf]. (ii) We use baseline CIFAR10 and CIFAR100 datasets [krizhevsky2009learning] and manually partition data over 50 federated learning users, with each holding 1,000 samples.
@@
(a) Severe Heterogeneity ($\alpha$=0.1)
(b) Moderate Heterogeneity ($\alpha$=0.5)
(c) Weak Heterogeneity ($\alpha$=1.0)
Figure 6: Visualization of training data distribution partitioned by the Dirichlet-based method. Best viewed in color.
Figure 5 visualizes the data distribution used for our evaluations (c.f., Sec 4). In this section, we employ two complementary dataset‐splitting strategies. First, in Sections 4.24.3, and 4.4, we use pathological (shard-based) partitioning, a common approach for simulating real-world data heterogeneity. Here, each class’s data is divided into fixed-size shards, and each client is assigned a subset of these shards without overlap. By varying both the number of shards per client and the shard size, we induce controlled quantity skew (different total data amounts) and label skew (uneven class distributions) across clients. In image-based tasks such as CIFAR10 and CIFAR100, this shard-based assignment also implicitly introduces feature-level skew, since different clients see different subsets of the underlying intra-class variations (e.g., orientation, background, and lighting), leading to mixed heterogeneity in both label composition and input characteristics.
Second, for all other experiments in Section 4, we adopt a Dirichlet-based dataset partitioning. Figure 6 shows the data distribution sampled via Dirichlet-based partitioning over the clients. Concretely, let $K$ be the number of classes and $C$ the number of clients. We first draw, for each client $i$, a class‐proportion vector
$\mathbf{p}^{(i)}\sim\mathrm{Dirichlet}(\alpha,\dots,\alpha)\,$,
where $\alpha>0$ is the concentration parameter. Then, for each class $k$, we allocate to client $i$ exactly
$n_{k}^{(i)}=\bigl\lfloor\,|\mathcal{D}_{k}|\;p_{k}^{(i)}\bigr\rfloor$
samples drawn without overlap from the global pool $\mathcal{D}_{k}$. In this way, each client ends up with a total of $\sum_{k}n_{k}^{(i)}$ samples whose class composition follows $\mathbf{p}^{(i)}$. By tuning $\alpha$, we control the degree of heterogeneity: (i) A small $\alpha$ (e.g. $\alpha=0.1$) yields highly skewed, non‐i.i.d. splits (clients see only a few classes); (ii) A moderate $\alpha$ (e.g. $\alpha=0.5$) induces moderate class imbalance; (iii) A large $\alpha$ (e.g. $\alpha\geq 1$) approaches the i.i.d. regime. This Dirichlet‐based method, therefore, provides a continuous spectrum of label and quantity skew, allowing us to systematically evaluate our algorithms under varying non‐i.i.d. conditions. Similar to the shard-based setting, when applied to image datasets, Dirichlet partitioning also induces mixed skew with clients differing not only in class proportions but also in the subset of visual factors they observe; resulting in both label and feature heterogeneity.
We note that these two widely adopted non-i.i.d. benchmarks (shard-based and Dirichlet-based partitioning) already create mixed heterogeneity patterns that combine label skew, quantity skew, and, in the case of image datasets, feature-level skew arising from intra-class variability.
Model. For each dataset, we utilized distinct model architectures, taking into account the specific challenges and characteristics inherent to each dataset. For FEMNIST, an MLP model featuring four linear layers was employed. We term this combination as “FEMNIST-DNN” in the rest of our evaluations. In the case of the CIFAR10 and CIFAR100 datasets, experiments were conducted using a two-layer CNN model [mcmahan2017communication] and a ResNet18 [he2016deep], respectively. These two cases are denoted as “CIFAR10-CNN” and “CIFAR100-ResNet18.” We note that the impact of temperature scaling during the local training process is a model/dataset agnostic phenomenon.
Layer Details Repetition FC 1 Linear(784, 512) ReLU() $\times 1$ FC 2 Linear(512, 256) ReLU() $\times 1$ FC 3 Linear(256, 128) ReLU() $\times 1$ FC 4 Linear(256, 62) $\times 1$
Layer Details Repetition Conv 1 Conv2d(3, 10, k=(5, 5), s=(1, 1)) ReLU() MaxPool2d(k=(2, 2)) $\times 1$ Conv 2 Conv2d(10, 20, k=(5, 5),s=(1, 1)) ReLU() MaxPool2d(k=(2, 2)) $\times 1$ FC 1 Linear(500, 256) ReLU() $\times 1$ FC 2 Linear(256, 10) $\times 1$
Table 1 and 2 show the detailed model architecture used for the experiments, respectively. For the ResNet18 model used to train the CIFAR100 dataset, we exploited PyTorch implementation [NEURIPS2019_9015] by only replacing the final classification head from Linear(512, 1000) to Linear(512, 100). It is important to note that low temperatures may cause unstable convergence in large models (e.g., ResNet, ViT). We recommend incorporating normalization layers to mitigate this effect. We further note that investigating methods to minimize instability caused by low temperatures with respect to different model architectures or components is an interesting direction for future work.
In real‐world deployments, clients often employ heterogeneous model architectures tailored to their specific applications and resource constraints. In this work, however, we restrict our evaluation to a homogeneous setting in which all clients share the same model architecture, thereby abstracting away variations in computational and communication capabilities. Extending our approach to support architectural heterogeneity and diverse client resources is an important avenue for future research, as it would more closely reflect practical federated learning scenarios.
Experiment Details. Each case underwent local training for 300 federated training rounds, utilizing FedAvg [mcmahan2017communication] as the baseline federated learning framework. FedAvg is a widely used framework for federated learning. In each round, the server selects 10 participating clients, employing Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.001 [ruder2016overview]. Each client trained its model for 10 local epochs, employing a batch size of 16, and utilized cross-entropy loss as the loss function. Unless specified otherwise, the experiment was conducted using the same hyperparameters as mentioned in this section. Table 3 shows the default hyperparameters used for the experiments. We note that we applied a learning rate decay of $10^{-5}$ to enhance training stability. Although the effect was minimal and did not alter the trend of the main results, we recommend its application.